OpenStudy (anonymous):

Find two consecutive integers whose sum is 67. Which of the following equations could be used to solve the problem? 2x = 67 2x + 1 = 67 2x + 2 = 67 x2 + 1= 67

OpenStudy (anonymous):

also Peggy had three times as many quarters as nickels. She had $1.60 in all. How many nickels and how many quarters did she have? Which of the following equations could be used to solve the problem? n + 3n = 1.60 n + 3n = 160 5n + 75n = 1.60 5n + 75n = 160

OpenStudy (anonymous):

also The cost to mail a package is $7 for the first 2 pounds and 30 cents for each additional ounce. Which of the following functions represents the cost to mail a package if x is the number of ounces over 2 pounds? f(x) = 7 + 0.3x f(x) = 14 + 0.3x f(x) = 0.3(x + 2)

OpenStudy (anonymous):

For the first one, I would say 2x+1=67 where x=33 so 2x+1 would equal (33+33)+1

OpenStudy (anonymous):

next What is the equation of the line that passes through (0, -2) and has a slope of 0? y = -2 y = 2 x = 2